OpenStudy (anonymous):

Could someone please help me on this, Let f(x) = [x^2+1]/[5+x] . Evaluate the following expression : f(a) - f(2) Thanks in advance :)

OpenStudy (anonymous):

\[f(a)=\frac{a^{2}+1}{5+a}...........f(2)=\frac{2^{2}+1}{5+2}=\frac{5}{7}\]

OpenStudy (anonymous):

\[f(a)-f(2) =\frac{7a^{2}+7-25-5a}{7(5+a)}=\frac{7a^{2}-5a-18}{7(a+5)}\]

OpenStudy (anonymous):

ohh i get now , thanks a lot for your help !